A ^{1}Σ^{+}−X̃^{1}Σ^{+} ELECTRONIC TRANSITION OF YOH IN THE VISIBLE REGION

K. ATHANASSENAS, C. T. KINGSTON, A. J. MERER, and \underline{J. R. D. PEERS}, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, B.C., Canada V6T 1Z1.; D. A. GILLETT, Lambda Physik, Inc., 3201 West Commercial Blvd., Fort Lauderdale, FL 33309..

Excitation spectra of the new molecules YOH and YOD have been recorded in the 500-625 nm wavelength region by laserinduced fluorescence, following reaction of laser-ablated yttrium with H_{2}O or D_{2}O under supersonic jet-cooled molecular beam conditions. Rotational analyses of ^{1}Σ^{+}−X̃^{1}Σ^{+} bands near 540 nm at high resolution gave the following substitution structure for the ground state:

X̃^{1}Σ^{+}: r(Y-O) = 1.948 Å; r(O-H) = 0.92_{1} Å.

Wavelength-resolved fluorescence spectra have given the ground state vibrational frequencies

YOH (YOD): ν_{2}, bend = 314 (247) cm^{−1}; ν_{3}, YO stretch = 674 (654) cm^{−1}

No evidence for the ν_{1} (OH stretch) mode was observed. Very complicated weak structure at the long wavelength end may possibly correspond to a singlet-triplet transition.